OpenStudy (kasper-kay):

assume no denomiator equals 0. simplify 15p^3q^2/ 18p^2q 3a^2b/9ab^2

OpenStudy (welshfella):

first find the GCF of 15 and 18 can you do that?

OpenStudy (kasper-kay):

yeah its 3

OpenStudy (welshfella):

so divide 15 and 18 by 3 to simplify the numerals

OpenStudy (welshfella):

now take the terms in p p^3 / p^2 = ?

OpenStudy (kasper-kay):

5p^3 / 6p^2 like that?

OpenStudy (welshfella):

5 and 6 is right but you need to work out p^3 / p^2 when you divide you subtract the esponents for example x^4 / x^2 = x(4-2) = x^2

OpenStudy (welshfella):

p^3 / p^2 = p^(3-2) = ?

OpenStudy (kasper-kay):

so p^1

OpenStudy (welshfella):

yes or just write p The 1 is understood now q^2 / q = ?

OpenStudy (welshfella):

q^2 / q^1 =

OpenStudy (kasper-kay):

so q

OpenStudy (welshfella):

= q^(2-1) = q

OpenStudy (welshfella):

so we get 5pq --- 6

OpenStudy (welshfella):

the second one is done in same way

OpenStudy (welshfella):